Coordinator for this week 15 : Martin Jäkel
Current PS / SPS User schedule v1.1.0
Problem with updating web page. Schedule attached to agenda.
Injector schedule (v1.1) for 2025 is available here: https://edms.cern.ch/document/3057281/latest
EuroLabs : We will soon send out the call for 2025. New day rates (tbc).
M.R. Jäkel, P. Martinengo
Proton Run 2025
Newly vacant beam slots :
Please contact SPS.Physics.Coordinator@cern.ch if you want to apply for any of theses slots.
Previous scheduled User looking for beam time :
Reminder: Beam time exceeding the limits of 2 weeks PS beam time and 1 week SPS beam time per year need the approval of one of these CERN committees: SPSC, LHCC, DRDC, INTC, RB or IEFC. Consider joining a DRD collaboration, if you require more beam time.
Target intensities:
See below or via ASM (accessible from within the CERN network).
Nothing to report.
Upcoming SPS MDs next weeks :
https://be-dep-ea.web.cern.ch/content/md-planning-north-area
Dedicated SPS MDs - if not anounced differently - are taking place from 8:00 to 18:00
Interventions in the ‘Transmission station area’ (close to the last collimator before EAR1), to mount a Diamond detector for additional beam monitoring.
Smooth data taking on:
• EAR1: C6D6 detector for the study of the 97,98Mo(n, g) reaction
• EAR2: study of the 87Sr(n, g) reaction
• NEAR: Study of radiation hardness for a diamond detector
Delivered protons
T09: Good operation. No issues to report. Magnets were degaussed.
T10: Good operation. Now running with 0.8s and maximum 7 GeV/c momentum. Many thanks to PS-OP colleagues for the quick setup and alignment onto the target.
T11: No user.
Smooth operation week in T8. Proton fluence fluence above the weekly reference value (>2.8E16 p/w). TID in CHARM about 530 Gy. Beam centering on T8 was excellent on both axes (>97% of the time center within +/- 2mm).
Cumulative plots available weekly on-line on the www.cern.ch/ps-irrad webpage (see for instance):
Plotting the beam transmission (see below the period 21/3 until today) at extraction in F61 (XSEC23/PS.BCT - black curve) and delivered to T8 (XSEC70/PS.BCT - red curve) it seems that the XSEC23 signal slowly drift over time. SY-BI colleagues already informed on Monday:
During the weekely acces of Wed. April 9th, new samples were added. In IRRAD, the experiments from SY-BI and samples from TE-MSC, EP-DT and LHCb TimeSPOT continue. In CHARM exchanged BE-CEM setups in the rack
Next access will be on Wednesday April 16th.
Taking beam data at low momentum, we have now switched to negative polarity.
We are seeing nice Cherenkov rings in the detector!
However at negative polarity we observed a 35 to 40 MeV/c offset in the momentum of muons and pions measured by our TOF compared to the beam momentum we expect for each run.
After degaussing the magnet, it reduced to about 25 to 30 MeV/c offset, but still seen in both muons and pions
We will continue scanning through the momenta up to 500 MeV/c and monitor this effect, and adapt our runs as needed.
Next week we have new trigger hardware arriving to enable us to take data at higher rates and higher momenta.
Finished data taking yesterday. A lot of different Timepix family detectors were tested from the last Friday (Timepix2, Timepix3 with different sensor materials, Timepix4). Data taking was smooth. Thanks a lot for the beam delivery and setting it up.
Taking data since yester evening, smooth installation
Beam quite nice, thank you for the fast setting up
General: different alignments through T4 target (the “wobbling”) measured in P42/H8/H6 together with a good optimum found.
H2/H4: Beam comissioning ongoing smoothly, information for possibilities of purities and rates has been communicated to the various users and the physics coordinator. H2-VLE to be comissioned on week 28.
H6: Beam commissioning going well, but CEDAR PMT replacements required a few accesses. H6 will be accessible on Saturday morning due to upstream beam tests, but otherwise in use.
H8: Beam commissioning going well, final checks ongoing for standard 180 GeV/c beam. XCETs to be checked.
M2: All beam modes checked for the beam commissioning. Some final checks planned for this week. Water pressure fault with Quad.065095 required access to BA81 on Wednesday.
P42/K12: Beam commissioning progressing well. P42 done and K12 nearly done, only focussing left at the CEDAR.
Next pre-comissioning plot for DPA : Week 23.
We are finalizing the commissioning of the system, we almost fixed a major problem with the GTK.
We are almost ready to use muon beam for detector calibrations.
Next Monday we planned to fill our CEDAR with H2 and the days after (2 days) will be dedicated to the commissionig of the CEDAR and then
we are ready to start with our physics program.
All set-ups almost ready
CMS-CSC
ATALS-MPI
EP-DT2
ECOGAS
CMS-RPC
USTC RPC
Straw Tracker parasitically in PPE134
We have started moving our setup into the area in order to save time during the installation phase.
Test beam period: 14th April – 30th April
Main goals:
Figure: An HGTD module mounted on the support frame.
DAQ setup: acquisition system will follow a configuration similar to the previous test beam based on EUDAQ2, MCP at 2500V as timing reference, Ethanol chiller for cold box cooling
Beam setup, same as in the previous test beams:
ISIEC submitted and safety check planned for Monday at 2PM.
Main goal: Characterization of irradiated Trenched Isolated Low Gain Avalanche Diodes (TI-LGADs), focusing on efficiency in the interpixel region. Several structures with varying gain layer-to-trench distance and trench widths will be tested.
Figure: Example of TI-LGADs structures to be testes
Planned Setup:
Figure: Previous setup, very similar to the planned one.
ISIEC submitted. Safety check planned for Tuesday at 11:00 AM.
Test of Scintillating Fibre mats for LHCb Upgrade II. We’re gonna test new fibre mats with and without mirror at the end and new SiPM readout.
Setup will be installed on the orange table in PPE168, like last years.
Will keep in touch with Maarten for beam characteristics.
ISIEC already submitted last Tuesday.
goals:
Due to the parasitic data-taking wrt GIF++, only basic technical tasks are considered
-test of the first DUNE tracker readout prototype
-test of the NA64 straw readout solution to obtain the best possible resolution
-continue common work with University of Michigan ATLAS sMDT team on ASD-based straw readout – test synchronization between ASD and reference tracker DAQs
Installation:
- relocation from H8 beam dump to PPE134 (crane): 11.04 15:00
- safety inspection: 14.04 09:00
Minutes by the respective speakers, edited by M. Jäkel, M. Malkin, and M. Schwinzerl